OpenStudy (anonymous):

please help with this one Solve the quantity x plus 5 divided by the quantity x squared plus x equals 1 divided by the quantity x squared plus x minus the quantity x minus 6 divided by the quantity x plus 1.

OpenStudy (blurbendy):

\[\frac{ (x + 5) }{ (x^2 + x) } = \frac{ 1 }{ x^2 + x - \frac{ x - 6 }{ x + 1 } }\]

OpenStudy (blurbendy):

that's what it looks like it's asking. would you agree?

OpenStudy (anonymous):



OpenStudy (blurbendy):

oh ok, i would start by getting everything on the left side, then we'll have to find a common denominator

OpenStudy (blurbendy):

x^2 + 1 is the least common denominator, so we need to get x + 1 to equal x^2 + 1 in order to do that, we'll have to multiply the top and bottom by something to get it to equal x^2 + 1

OpenStudy (blurbendy):

any ideas?

OpenStudy (anonymous):

hmm im not sure

OpenStudy (blurbendy):

x* (x + 1) = x^2 + x, so you need to multiply the top and bottom of (x - 6) / (x +1) by x

OpenStudy (blurbendy):

understand?